Grade 12Algebra

Find the product, using suitable properties:
7 × (50 – 2)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

The given equation is in the form of Distributive law of Multiplication over Subtraction.
= a × (b - c) = (a × b) - (a × c)
= (7 × 50)–(7 × 2)
= 350–14
= 336

Thanks